OpenStudy (anonymous):

length of arc AB y = degree of angle r = radius r=5 y=69 answer rounded to two decimal places 2.53?

OpenStudy (anonymous):

69 degree= 69*pi/180 radian length of arc= radius*subtended angle at centre in radian= 5*69*pi/180=6.02 approx

OpenStudy (anonymous):

its a circle sorry forgot to mention that so its 69/360

OpenStudy (callisto):

Arc = 2(pi)(r) * angle in degree / 360 = 2 pi (5) (69)/360 = 6.02

OpenStudy (anonymous):

how do u get 69/360, @suzy?

OpenStudy (anonymous):

thats the angle of the circle 69/ 360

OpenStudy (anonymous):



OpenStudy (callisto):